2. The rate of the reaction A + 3B + C + 2D was reported as 2.7 mol. dm-3.5-1. a. State the rates of formation (and consumpti can you answer part b and part c?
Answer #1

A 3B- C-12D rate : 2.7 moldminis- = e râle of formetin of a Thold (a) rate of reachioa (r) = -102 - d[0) = 2.7 moldn-us=]] -